Solve the integral of tanhxsech
1
Expert's answer
2021-07-22T08:41:55-0400
"\int \tanh x\text{sech} xdx=\int\dfrac{\sinh x}{\cosh^2 x}dx"

"u=\cosh x, d=\sinh xdx"


"\int \tanh x\text{sech} xdx=\int\dfrac{\sinh x}{\cosh^2 x}dx"

"=\int\dfrac{du}{ u^2}=-\dfrac{1}{u}+C=-\dfrac{1}{\cosh x}+C="

"=-\text{sech} x"